1/3(X-4)+ 5/6(X+2)= X+3

Thanks.

Answer by stanbon(75887)

1/3(X-4)+ 5/6(X+2)= X+3
---------
Multiply thru by 6 to get rid of those fractions:
2(x-4) + 5(x+2) = 6(x+3)
2x-8+5x+10 = 6x+18
x = 16
===========
Cheers,
Stan H.